# Principal Component Copulas for Capital Modeling and Systemic Risk

### v2

@@ Abstract
We introduce a class of copulas that we call Principal Component Copulas (PCCs). This class combines the strong points of copula-based techniques with principal component-based models, which results in flexibility when modelling tail dependence along the most important directions in high-dimensional data. We obtain theoretical results for PCCs that are important for practical applications. In particular, we derive tractable expressions for the high-dimensional copula density, which can be represented in terms of characteristic functions. We also develop algorithms to perform Maximum Likelihood and Generalized Method of Moment estimation in high-dimensions and show very good performance in simulation experiments. Finally, we apply the copula to the international stock market in order to study systemic risk. We find that PCCs lead to excellent performance on measures of systemic risk due to their ability to distinguish between parallel market movements, which increase systemic risk, and orthogonal movements, which reduce systemic risk. As a result, we consider the PCC promising for internal capital models, which financial institutions use to protect themselves against systemic risk.
@@ 1 Introduction
Modelling the dependence structure of a multivariate distribution plays a crucial role in finance and risk management. Copulas are a popular approach, since they allow to separate the choices for the marginal distributions from the dependence structure [1, 2]. The Gaussian copula has popularized the use of copulas in finance due to its simplicity, especially in high-dimensional applications [3]. However, its disadvantage is that it cannot capture tail dependence. As a consequence, the Student t copula and its extensions have become popular [4]. Although these copulas are often used in practice, also other copula methods have become available that enhance flexibility. Vine copulas are constructed by sequentially applying bivariate copulas to build higher-dimensional copulas [5]. The resulting pair copula constructions are flexible, but model complexity can quickly increase, which is a disadvantage for internal capital models of financial institutions.
More recently, factor copulas have received a steady increase of interest in the literature [6, 7, 8, 9, 10]. Oh and Patton have applied factor copulas to stock returns and have shown an enhanced ability in capturing systemic risk [8]. Duan et al. find good performance for Value-at-Risk forecasts using factor copulas [10]. As a result, such copulas are promising for capital modelling. Although the attention for factor copulas has been increasing, there also remain challenges. In general, no analytic expressions exist for the copula density, which makes estimation nontrivial. Oh and Patton have developed a simulation-based estimation technique, but this method is numerically challenging, which has limited the number of estimated parameters [8]. Another important challenge for factor copulas is to determine the appropriate factor structure. For return data, the factors are often based on sectoral codes [7, 8, 9]. However, this approach is suboptimal compared to a data-driven clustering approach [11]. This shows that finding the appropriate factor structure is in general a non-trivial task.
@@
Another contribution of our article is that we develop efficient algorithms to estimate high-dimensional PCCs. We propose the Generalized Hyperbolic (GH) distribution as stochastic generator family, because of its flexibility and its analytic tractability. In order to achieve Maximum Likelihood Estimation (MLE), we develop numerical methods to accurately determine the high-dimensional copula density at limited computational costs. We propose the Fast Fourier Transform [15] and the COS method [16] to efficiently compute the required one-dimensional integrals over characteristic functions. Due to the rapid evaluation of the copula density, we achieve efficient MLE for flexible copula structures. In high dimensions, the number of correlation parameters grows rapidly. Full MLE remains feasible only when additional structure is applied to the correlation matrix, so that it can be parameterized in a parsimonious way. Although such restricted models can be beneficial, it is usually not desirable to make strong a priori assumptions on the correlation matrix. Therefore, we also propose a second estimation scheme. For high-dimensional copulas, it is natural to make a distinction between correlation parameters (second-moment parameters) and shape parameters (higher-moment parameters). We develop a novel algorithm that is designed to efficiently estimate a large number of correlation parameters using the method of moments and a small number of shape parameters based on MLE. We show excellent performance of this algorithm in high-dimensional simulation experiments.
Our final contribution is that we apply the developed copula to the international stock market in order to study global diversification and systemic risk. Recently, several copula studies have appeared to analyze systemic risk in bond markets [17], stock markets [8], credit derivative markets [18] and commodity markets [19]. We expand upon existing studies by introducing the market distress cube to empirically study systemic risk in three directions, namely more extreme crashes, a larger global market size, and an increasing market share. We also develop a binomial test to assess the ability of various copula specifications to capture high-dimensional dependence. We find that PCCs outperform existing copulas based on information criteria and measures of systemic risk. The enhanced performance arises from their ability to distinguish between parallel variation, which increases systemic risk and reduces diversification, and orthogonal variation, which reduces systemic risk and enhances diversification. As a result, we consider the PCC promising for quantitative risk management and capital modeling, which financial institutions use to protect themselves against systemic risk.
@@ 2 Principal Component Copulas
The main advantage of a copula is the possibility to separate the modelling of marginal distributions from the modelling of the dependence structure [1, 2]. In high dimensions, implicit copulas are commonly applied, which are constructed from analytically tractable multivariate distributions. Well-known examples of implicit copulas are Gaussian copulas, normal mixture copulas such as the (skew) $t$ copula, and factor copulas. A recent review of implicit copulas is given by Smith [20]. In general, the joint distribution function $F_{\bm{Y}}(y_{1},...,y_{d})$ for the stochastic vector $\mathbf{Y}$ can be expressed as $F_{\mathbf{Y}}(y_{1},...,y_{d})=C_{\mathbf{Y}}(F_{Y_{1}}(y_{1}),...,F_{Y_{d}}(y_{d}))$ with $F_{Y_{i}}$ the marginal distribution functions and $C_{\mathbf{Y}}(u_{1},...,u_{d})$ the implicit copula belonging to $F_{\bm{Y}}$. If the marginal distributions are continuous, the implicit copula is unique, which is the setting of relevance to us. The copula density is then given by:
@@ 2.1 Data generating process for the implicit copula
The copula construction therefore starts from uncorrelated generators or PCs. The correlation structure of the implicit copula is consequently determined by the diagonal covariance matrix $\bm{\Lambda}$ for the generators and by the orthogonal transformation matrix $\bm{W}$. Flexible tail dependence structures for the copula can be obtained using appropriate families of generator distributions. If we specify the Data Generating Process (DGP) for the stochastic vector ${\bm{P}}$ by its continuous multivariate distribution function $F_{\bm{P}}$, the distribution function $F_{\bm{Y}}$ and the corresponding copula $C_{\bm{Y}}$ are obtained by the linear transformation of Eq. (2). We therefore arrive at the following definition.
@@ 2.2.1 Hyperbolic-Normal PCC
$f_{\rm HB}(x)=\frac{\gamma}{2\alpha\delta K_{1}(\delta\gamma)}e^{-\alpha\sqrt{\delta^{2}+(x-\mu)^{2}}+\beta(x-\mu)}.$ (3)
Here, $\gamma=\sqrt{\alpha^{2}-\beta^{2}}$ with $\alpha>\beta$ and $K_{\lambda}$ is a modified Bessel function of the second kind, $\mu$ is a location parameter, $\delta$ is a scale parameter, $\alpha+\beta$ describes the lower tail, and $\alpha-\beta$ the upper tail. The expressions for the mean and variance of the hyperbolic distribution are known analytically. The parameter $\mu$ is chosen such that the mean is zero. The variance $\Lambda$ is a monotonically increasing function of $\delta$, and we denote the corresponding hyperbolic distribution as ${\rm HB}(\alpha,\beta,\Lambda)$.
@@
We start by considering the two-dimensional case for simplicity. Then, the second PC is normally distributed as $P_{2}\sim N(0,\Lambda_{2})$ with $\Lambda_{2}=d-\Lambda_{1}$. Suppose there is a positive linear correlation $\rho$ between the risk factors $Y_{i}$, such that $10<\rho<$ 1, and we apply PCA to the correlation matrix of $\mathbf{Y}$. The first PC weight vector is then the parallel movement of the risk variables $\mathbf{w}_{1}=[1,1]^{T}/\sqrt{2}$, while the second PC weight vector is the anti-parallel movement $\mathbf{w}_{2}=[1,-1]^{T}/\sqrt{2}$. A negative correlation would cause the anti-parallel vector to be the first PC. Moreover, we have $\Lambda_{2}=d-\Lambda_{1}$ with $d=2$. This shows we can parametrize the implicit copula by either $\rho$ or $\Lambda_{1}$. We prove the following proposition for the tail dependence of the two-dimensional HB-N PCC.
@@
We start by noting that the lower tail of the hyberbolic density is governed by: $f_{P_{1}}(x)\propto e^{(\alpha_{1}+\beta_{1})x}$. We proceed by evaluating the expression for the (lower) tail dependence coefficient $\eta_{l}$ (see e.g. [2] or [22]):
@@
In the third step, the leading contribution of the integrals arises from the regions centered around $y_{q}$ and $\sqrt{2}y_{q}$, respectively. The last step consists of evaluating two standard integrals. ∎
@@
For an elliptical distribution, tail dependence is generated when the density generator is regularly varying [13]. For a PCC, we have shown explicitly that we can also generate tail dependence with exponentially and Gaussian tailed variables, since there is no elliptical symmetry assumed.
@@ 2.2.2 Skew $t_{1}$-$t_{d-1}$ PCC
More general PCC constructions are obtained with the generalized hyperbolic distribution (GH) as generator. The GH distribution has a stochastic representation as a normal mean-variance mixture with the generalized inverse Gaussian (GIG) distribution as the stochastic mixing variable [22]. The corresponding class of implicit copulas is called a normal mixture copula. We will consider PCCs with multiple stochastic mixing factors, which will result in convolutions of the GH distribution. We start with the density for the $d$-dimensional GH distribution [23]:
$f_{\rm GH}(\bm{x})=\frac{(\psi/\chi)^{\lambda/2}\alpha^{d/2-\lambda}}{(2\pi)^{d/2}|\bm{\Sigma}|^{1/2}K_{\lambda}(\sqrt{\chi\psi})}\frac{e^{\bm{\beta}(\bm{x}-\bm{\mu})}K_{\lambda-d/2}\left(\alpha\sqrt{\chi+(\bm{x}-\bm{\mu})^{\prime}\bm{\Sigma}^{-1}(\bm{x}-\bm{\mu})}\right)}{\left(\sqrt{\chi+(\bm{x}-\bm{\mu})^{\prime}\bm{\Sigma}^{-1}(\bm{x}-\bm{\mu})}\right)^{d/2-\lambda}}$ (7)
Here, we have introduced the shorthand notations $\bm{\beta}=\bm{\Sigma}^{-1}\bm{\gamma}$ and $\alpha=\sqrt{\psi+\bm{\gamma}^{\prime}\bm{\Sigma}^{-1}\bm{\gamma}}$. Importantly, also the characteristic function for the GH distribution is analytically known, which is given by [23]:
$\phi_{\rm GH}(\bm{t})=\frac{e^{i\bm{t}^{\prime}\bm{\mu}}\psi^{\lambda/2}}{(\psi+\bm{t}^{\prime}\bm{\Sigma}\bm{t}/2-i\bm{\beta}^{\prime}\bm{t})^{\lambda/2}}\frac{K_{\lambda}\left(\sqrt{\chi(\psi+\bm{t}^{\prime}\bm{\Sigma}\bm{t}/2-i\bm{\beta}^{\prime}\bm{t})}\right)}{K_{\lambda}(\sqrt{\chi\psi})}.$ (8)
Many relevant distributions in finance are part of the GH family. When $\lambda=-1/2$, the normal-inverse Gaussian (NIG) distribution is obtained, while $\lambda=1$ results in the hyperbolic distribution. When $\lambda=-\nu/2$, $\chi=\nu$, and $\psi=0$, the multivariate skew $t$ distribution is obtained. The corresponding implicit copula is known as the GH skew $t$ copula. It has been used in several high-dimensional financial applications [4, 17, 9, 11]. When the skewness parameter vector $\bm{\gamma}$ goes to zero, the multivariate $t$ distribution and the corresponding $t$ copula with degrees-of-freedom (DoF) parameter $\nu$ are retrieved [4]. This shows the flexibility of the GH family, resulting in a convenient generator.
@@
$P_{j}=\begin{cases}\mu_{1}+\gamma_{1}V+\Sigma_{1,1}\sqrt{V}Z_{1},&\text{ for }j=1\\ \Sigma_{j,j}\sqrt{W}Z_{j},&\text{ for }j>1\end{cases}$ (9)
@@
Here, $Z_{j}\sim N(0,1)$, $V\sim{\rm IG}(\nu_{1}/2,\nu_{1}/2)$ and $W\sim{\rm IG}(\nu_{>1}/2,\nu_{>1}/2)$ are independent stochastic variables with $\nu_{1}$ the DoF parameter for the first generator and $\nu_{>1}$ the parameter for the higher generators. As a result, the generators $P_{j}$ are uncorrelated, but the higher PCs with $j>1$ are dependent due to the common mixing variable $W$. Since the expectation value and the variance of the GH distribution have analytic expressions [23], we can determine $\mu_{1}$ and $\Sigma_{j,j}$ such that for all $j$ the mean of $P_{j}$ is zero and the variance of $P_{j}$ equals $\Lambda_{j}$. For the GH skew $t$ distribution, these analytic expressions are particularly simple [4], so that we obtain:
$\mu_{1}=-\frac{\nu_{1}\gamma_{1}}{(\nu_{1}-2)},\quad\Sigma_{1,1}=\frac{(\nu_{1}-2)}{\nu_{1}}\Lambda_{1}-\frac{2\nu_{1}}{(\nu_{1}-2)^{2}(\nu_{1}-4)}\gamma_{1}^{2}.$ (10)
@@
Similarly, we have $\Sigma_{j,j}=(\nu_{>1}-2)\Lambda_{j}/\nu_{>1}$ for the higher generators with $j>1$. We require $\nu_{1}>4$ and $\nu_{>1}>2$ for the variance $\Lambda_{j}$ to remain finite. This is because the heaviest tail of the GH Skew $t$ distribution decays proportional to $|x|^{-\nu/2-1}$ [24], while the $t$ distribution decays proportional to $|x|^{-\nu-1}$.
@@
The first copula generator $P_{1}$ is used to generate asymmetric tail dependence. The higher PCs do not incorporate skewness. The Skew $t_{1}$-$t_{d-1}$ PCC is then based on the following specification:
$P_{1}\sim{\rm Skew}\,t_{1}(\nu_{1},\Lambda_{1},\gamma_{1}),\quad\bm{P}_{>1}\sim t_{d-1}(\nu_{>1},\bm{\Lambda}_{>1}).$ (11)
@@
Here, Skew $t_{1}(\nu_{1},\Lambda_{1},\gamma_{1})$ denotes the one-dimensional GH skew $t$ distribution with zero mean, variance $\Lambda_{1}$, degrees-of-freedom parameter $\nu_{1}$ and skewness parameter $\gamma_{1}$, while $t_{d-1}(\nu_{>1},\bm{\Lambda}_{>1})$ denotes the ($d-1$)-dimensional $t$ distribution with diagonal covariance matrix $\bm{\Lambda}_{>1}$. The matrix $\bm{\Lambda}_{>1}$ is obtained from $\bm{\Lambda}$ by omitting the first row and column containing $\Lambda_{1}$. The DGP for $\bm{Y}$ follows from the linear transformation $\bm{Y}=\bm{W}\bm{P}$. The pairwise tail dependence parameters, that follow from the convolution of two independent factors with regularly varying tails, have been analytically obtained by several authors [8, 14, 25]. Their results show that nontrivial tail dependence structures are obtained from the convolution of heavy-tailed factors with a common tail index. We conclude that the proposed implicit copula has the following combination of favorable properties specific to PCCs. It is skewed along the first generator, which is the main common factor. The distribution is elliptical in the subspace of the higher generators, which are the orthogonal directions to the main factor resulting in diversification. The stochastic structure is based on two stochastic mixing variables. This differs from the commonly applied (skew) $t$ copula, which has a single stochastic mixing factor $W$. In that case, the stochastic variance factor $W$ grows for all directions simultaneously in extreme scenarios. The variation orthogonal to the parallel movement is not a priori expected to grow during a crash, because this orthogonal variation enhances diversification. Consequently, we find in our empirical study of Section 4 that the Skew $t_{1}$-$t_{d-1}$ PCC specification leads to improvements in terms of information criteria and in capturing systemic risk compared to other copulas.

### v3

@@ Abstract
We introduce a class of copulas that we call Principal Component Copulas (PCCs). This class combines the strong points of copula-based techniques with principal component analysis (PCA), which results in flexibility when modelling tail dependence along the most important directions in high-dimensional data. We obtain theoretical results for PCCs that are important for practical applications. In particular, we derive tractable expressions for the high-dimensional copula density, which can be represented in terms of characteristic functions. We also develop algorithms to perform Maximum Likelihood and Generalized Method of Moment estimation in high-dimensions and show very good performance in simulation experiments. Finally, we apply the copula to the international stock market to study systemic risk. We find that PCCs lead to excellent performance on measures of systemic risk due to their ability to distinguish between parallel and orthogonal movements in the global market, which have a different impact on systemic risk and diversification. As a result, we consider the PCC promising for capital models, which financial institutions use to protect themselves against systemic risk.
@@ 1 Introduction
Modelling the dependence structure of a multivariate distribution plays a crucial role in finance and risk management. Copulas are a popular approach, since they allow to separate the choices for the marginal distributions from the dependence structure [1, 2]. The Gaussian copula has popularized the use of copulas in finance due to its simplicity, especially in high-dimensional applications [3]. However, its disadvantage is that it cannot capture tail dependence. As a consequence, the Student t copula and its extensions have become popular [4]. Although these copulas are often used in practice, also other copula methods have become available that enhance flexibility. Vine copulas are constructed by sequentially applying bivariate copulas to build higher-dimensional copulas [5]. The resulting pair copula constructions are flexible, but model complexity can quickly increase, which can be a disadvantage for internal capital models of financial institutions.
More recently, factor copulas have received a steady increase of interest in the literature [6, 7, 8, 9, 10]. Factor copulas for stock returns have shown good performance in capturing systemic risk [8] and in Value-at-Risk forecasts [10]. As a result, such copulas are promising for capital modelling. Although the attention for factor copulas has been increasing, there also remain challenges. In general, no analytic expressions exist for the copula density, which makes estimation nontrivial. [8] have developed a simulation-based estimation technique, but this method is numerically challenging, which has limited the number of estimated parameters. Another important challenge for factor copulas is to determine the appropriate factor structure. For return data, the factors are often based on sectoral codes [7, 8, 9]. However, this approach is suboptimal compared to a data-driven clustering approach [11]. This shows that finding the appropriate factor structure is in general a non-trivial task.
@@
Another contribution of our article is that we develop efficient algorithms to estimate high-dimensional PCCs. We propose the Generalized Hyperbolic (GH) distribution as stochastic generator family, because of its flexibility and its analytic tractability. In order to achieve Maximum Likelihood Estimation (MLE), we develop numerical methods to accurately determine the high-dimensional copula density at limited computational costs. We propose the Fast Fourier Transform [15] and the COS method [16] to efficiently compute the required one-dimensional integrals over characteristic functions. Due to the fast evaluation of the copula density, we achieve efficient MLE for flexible copula structures. In high dimensions, the number of correlation parameters grows rapidly. Full MLE remains feasible only when additional structure is applied to the correlation matrix, so that it can be parameterized in a parsimonious way. Although such restricted models can be beneficial, it is usually not desirable to make strong a priori assumptions on the correlation matrix. Therefore, we also propose a second estimation scheme. For high-dimensional copulas, it is natural to make a distinction between correlation parameters (second-moment parameters) and shape parameters (higher-moment parameters). We develop a novel algorithm that is designed to efficiently estimate a large number of correlation parameters using the method of moments and a small number of shape parameters based on MLE. We show excellent performance of this algorithm in high-dimensional simulation experiments.
Our final contribution is that we apply the developed copula to the international stock market in order to study global diversification and systemic risk. Recently, several copula studies have appeared to analyze systemic risk in bond markets [17], stock markets [8], credit derivative markets [18] and commodity markets [19]. We expand upon existing studies by introducing the market distress cube to empirically study systemic risk in three directions, namely more extreme crashes, a larger global market size, and an increasing market share. We also develop a binomial test to assess the ability of various copula specifications to capture high-dimensional dependence. We find that PCCs outperform existing copulas based on information criteria and measures of systemic risk. The enhanced performance arises from their ability to distinguish between parallel movements and orthogonal movements in the global market, which have different consequences for systemic risk and diversification. As a result, we consider the PCC promising for quantitative risk management and capital modeling, which financial institutions use to protect themselves against systemic risk.
@@ 2 Principal Component Copulas
The main advantage of a copula is the possibility to separate the modelling of marginal distributions from the modelling of the dependence structure [1, 2]. In high dimensions, implicit copulas are commonly applied, which are constructed from analytically tractable multivariate distributions. Well-known examples of implicit copulas are Gaussian copulas, normal mixture copulas such as the (skew) $t$ copula, and factor copulas. A recent review of implicit copulas is given by [20]. In general, the joint distribution function $F_{\bm{Y}}(y_{1},...,y_{d})$ for the stochastic vector $\mathbf{Y}$ can be expressed as $F_{\mathbf{Y}}(y_{1},...,y_{d})=C_{\mathbf{Y}}(F_{Y_{1}}(y_{1}),...,F_{Y_{d}}(y_{d}))$ with $F_{Y_{i}}$ the marginal distribution function for index $i$ and $C_{\mathbf{Y}}(u_{1},...,u_{d})$ the implicit copula belonging to $F_{\bm{Y}}$. If the marginal distributions are continuous, the implicit copula is unique, which is the setting of relevance to us. The copula density is then given by:
@@ 2.1 Data generating process for the implicit copula
The copula construction therefore starts from uncorrelated generators or PCs. The correlation structure of the implicit copula is consequently determined by the diagonal covariance matrix $\bm{\Lambda}$ for the generators and by the orthogonal transformation matrix $\bm{W}$. Flexible tail dependence structures for the copula can be obtained using appropriate families of generator distributions. If we specify the Data Generating Process (DGP) for the stochastic vector ${\bm{P}}$ by its continuous multivariate distribution function $F_{\bm{P}}$, then the distribution function $F_{\bm{Y}}$ and the corresponding copula $C_{\bm{Y}}$ are obtained by the linear transformation of Eq. (2). We therefore arrive at the following definition.
@@ 2.2.1 Hyperbolic-Normal PCC
$f_{\rm HB}(x)=\frac{\sqrt{\psi/\chi}}{2\alpha K_{1}(\sqrt{\chi\psi})}e^{-\alpha\sqrt{\chi+(x-\mu)^{2}}+\beta(x-\mu)}.$ (3)
Here, $\psi=\alpha^{2}-\beta^{2}$ with $\alpha>\beta$ and $K_{\lambda}$ is a modified Bessel function of the second kind, $\mu$ is a location parameter, $\chi$ is a scale parameter, $\alpha+\beta$ describes the lower tail, and $\alpha-\beta$ the upper tail. The expressions for the mean and variance of the hyperbolic distribution are known analytically. The parameter $\mu$ is chosen such that the mean is zero. The variance $\Lambda$ is a monotonically increasing function of $\chi$. We denote the corresponding hyperbolic distribution with mean zero and variance $\Lambda$ as ${\rm HB}(\alpha,\beta,\Lambda)$.
@@
We start by considering the two-dimensional case for simplicity. Then, the second PC is normally distributed as $P_{2}\sim N(0,\Lambda_{2})$ with $\Lambda_{2}=d-\Lambda_{1}$. Suppose there is a positive linear correlation $\rho$ between the risk factors $Y_{i}$, such that $10<\rho<$ 1, and we apply PCA to the correlation matrix of $\mathbf{Y}$. The first PC weight vector is then the parallel movement of the risk variables $\mathbf{w}_{1}=[1,1]^{T}/\sqrt{2}$, while the second PC weight vector is the anti-parallel movement $\mathbf{w}_{2}=[1,-1]^{T}/\sqrt{2}$. A negative correlation would cause the anti-parallel vector to be the first PC. Moreover, we have $\Lambda_{2}=d-\Lambda_{1}$ with $d=2$. This means we can parametrize the implicit copula by either $\rho$ or $\Lambda_{1}$. A contour plot of the Hyperbolic-Normal copula density with Gaussian marginal distributions is given in Figure 1. We prove the following proposition for the tail dependence of the two-dimensional HB-N PCC.
@@
We start by noting that the lower tail of the hyberbolic density is governed by: $f_{P_{1}}(x)\propto e^{(\alpha_{1}+\beta_{1})x}$. We proceed by evaluating the expression for the (lower) tail dependence coefficient $\eta_{l}$ [22, 2]:
@@
In the third step, the leading contribution of the integrals arises from the regions centered around $y_{q}$ and $\sqrt{2}y_{q}$, respectively. The two integral expressions ($I_{1}$ and $I_{2}$) are standard integrals, for which we obtain:
$I_{1}$ $=$ $2\sqrt{\pi}\sigma_{2}e^{\sqrt{2}(\alpha+\beta)y_{q}+(\alpha+\beta)^{2}\sigma_{2}^{2}/2}\Phi(-(\alpha+\beta)\sigma_{2})$
$I_{2}$ $=$ $\sqrt{2}\sqrt{\pi}\sigma_{2}e^{\sqrt{2}(\alpha+\beta)y_{q}+(\alpha+\beta)^{2}\sigma_{2}^{2}/2}$ (7)
We finalize by inserting the expressions from Eq. (2.2.1) into Eq. (6). ∎
@@
For an elliptical distribution, tail dependence is generated when the random vector is regularly varying [13]. For a PCC, we have shown explicitly that we can also generate tail dependence with exponentially and Gaussian tailed variables, since there is no elliptical symmetry assumed.
@@ 2.2.2 Skew $t_{1}$-$t_{d-1}$ PCC
More general PCC constructions are obtained with the generalized hyperbolic distribution (GH) as generator. The GH distribution has a stochastic representation as a normal mean-variance mixture with the generalized inverse Gaussian (GIG) distribution as the stochastic mixing variable [22]. The corresponding class of implicit copulas is called a normal mixture copula. We will consider PCCs with multiple stochastic mixing factors, which will result in convolutions of the GH distribution. We start with the density for the $d$-dimensional GH distribution, for which various parameterizations exist [23]. We use:
$f_{\rm GH}(\bm{x})=\frac{(\psi/\chi)^{\lambda/2}\alpha^{d/2-\lambda}e^{(\bm{x}-\bm{\mu})\bm{\beta}}}{(2\pi)^{d/2}|\bm{\Sigma}|^{1/2}K_{\lambda}(\sqrt{\chi\psi})}\frac{K_{\lambda-d/2}\left(\alpha\sqrt{\chi+(\bm{x}-\bm{\mu})^{\prime}\bm{\Sigma}^{-1}(\bm{x}-\bm{\mu})}\right)}{\left(\sqrt{\chi+(\bm{x}-\bm{\mu})^{\prime}\bm{\Sigma}^{-1}(\bm{x}-\bm{\mu})}\right)^{d/2-\lambda}}$
Here, we have introduced the notations $\bm{\beta}=\bm{\Sigma}^{-1}\bm{\gamma}$ and $\alpha=\sqrt{\psi+\bm{\gamma}^{\prime}\bm{\Sigma}^{-1}\bm{\gamma}}$, while $\lambda\in\mathbb{R}$, $\chi>0$, $\psi>0$, $\bm{\mu}\in\mathbb{R}^{d}$, $\bm{\gamma}\in\mathbb{R}^{d}$ and $\bm{\Sigma}\in\mathbb{R}^{d\times d}$ a positive definite matrix. Importantly, also the characteristic function for the GH distribution is analytically known [23], which is given by:
$\phi_{\rm GH}(\bm{t})=\frac{e^{i\bm{t}^{\prime}\bm{\mu}}\psi^{\lambda/2}}{(\psi+\bm{t}^{\prime}\bm{\Sigma}\bm{t}-2i\bm{t}^{\prime}\bm{\gamma})^{\lambda/2}}\frac{K_{\lambda}\left(\sqrt{\chi(\psi+\bm{t}^{\prime}\bm{\Sigma}\bm{t}-2i\bm{t}^{\prime}\bm{\gamma})}\right)}{K_{\lambda}(\sqrt{\chi\psi})}.$ (8)
Many relevant distributions in finance are part of the GH family, since $\lambda=-1/2$ results in the normal-inverse Gaussian (NIG) distribution, while $\lambda=(d+1)/2$ results in the hyperbolic distribution. When $\lambda=-\nu/2$, $\chi=\nu$ and $\psi\rightarrow 0$, the multivariate skew $t$ distribution is obtained. The corresponding implicit copula is known as the GH skew $t$ copula, which is frequently used in financial applications [4, 17, 9, 11]. When the skewness parameter vector $\bm{\gamma}$ goes to zero, the multivariate $t$ distribution and the $t$ copula with degrees-of-freedom (DoF) parameter $\nu$ are retrieved [4].
Figure 1: Contour plots for bivariate copula densities generated by the Hyperbolic-Normal and the Skew $t_{1}$-$t_{1}$ Principal Component Copulas. For the visualization, we choose for the standard normal distribution as marginal distribution, so that the Gaussian copula results in elliptical contours. Because the PCC examples have a skewed and leptokurtic distribution along the first Principal Component, enhanced probabilities of joint downward movements are generated.
@@
$P_{j}=\begin{cases}\mu_{1}+\gamma_{1}V+\sigma_{1}\sqrt{V}Z_{1},&\text{ for }j=1\\ \sigma_{j}\sqrt{W}Z_{j},&\text{ for }j>1\end{cases}$ (9)
@@
Here, $Z_{j}\sim N(0,1)$, $V\sim{\rm IG}(\nu_{1}/2,\nu_{1}/2)$ and $W\sim{\rm IG}(\nu_{>}/2,\nu_{>}/2)$ are independent stochastic variables, while $\sigma_{j}=\sqrt{\Sigma_{j,j}}$, $\nu_{1}$ is the DoF parameter for the first generator and $\nu_{>}$ the parameter for the higher generators. As a result, the generators $P_{j}$ are uncorrelated, but the higher PCs with $j>1$ are dependent due to the common mixing variable $W$. Since the expectation value and the variance of the GH distribution have analytic expressions [23], we can determine $\mu_{1}$ and $\Sigma_{j,j}$ such that for all $j$ the mean of $P_{j}$ is zero and the variance of $P_{j}$ equals $\Lambda_{j}$. For the GH skew $t$ distribution, these analytic expressions are particularly simple [4], so that we obtain:
$\mu_{1}=-\frac{\nu_{1}\gamma_{1}}{(\nu_{1}-2)},\quad\Sigma_{1,1}=\frac{(\nu_{1}-2)}{\nu_{1}}\Lambda_{1}-\frac{2\nu_{1}}{(\nu_{1}-2)(\nu_{1}-4)}\gamma_{1}^{2}.$ (10)
@@
Similarly, we have $\Sigma_{j,j}=(\nu_{>}-2)\Lambda_{j}/\nu_{>}$ for the higher generators with $j>1$. We require $\nu_{1}>4$ and $\nu_{>}>2$ for the variance $\Lambda_{j}$ to remain finite. This is because the heaviest tail of the skew $t$ distribution decays as $|x|^{-\nu/2-1}$ [24], while the $t$ distribution decays as $|x|^{-\nu-1}$.
@@
The first copula generator, $P_{1}$, is used to generate asymmetric tail dependence. The higher PCs, $\bm{P}_{>}$, do not incorporate skewness. The Skew $t_{1}$-$t_{d-1}$ PCC is then based on the following specification:
$P_{1}\sim{\rm Skew}\,t_{1}(\nu_{1},\Lambda_{1},\gamma_{1}),\quad\bm{P}_{>}\sim t_{d-1}(\nu_{>},\bm{\Lambda}_{>}).$ (11)
@@
Here, Skew $t_{1}(\nu_{1},\Lambda_{1},\gamma_{1})$ denotes the one-dimensional GH skew $t$ distribution with zero mean, variance $\Lambda_{1}$, degrees-of-freedom parameter $\nu_{1}$ and skewness parameter $\gamma_{1}$, while $t_{d-1}(\nu_{>},\bm{\Lambda}_{>})$ denotes the ($d-1$)-dimensional $t$ distribution with diagonal covariance matrix $\bm{\Lambda}_{>}$. The matrix $\bm{\Lambda}_{>}$ is obtained from $\bm{\Lambda}$ by omitting the first row and column containing $\Lambda_{1}$. The DGP for $\bm{Y}$ follows from the linear transformation $\bm{Y}=\bm{W}\bm{P}$. A contour plot of the Skew $t_{1}$-$t_{d-1}$ copula density with Gaussian marginal distributions is given in Figure 1. The pairwise tail dependence parameters, that follow from the convolution of two independent factors with regularly varying tails, have been analytically obtained by several authors [25, 14, 8]. Their results show that nontrivial tail dependence structures are obtained from the convolution of heavy-tailed factors with a common tail index.
We conclude that the proposed implicit copula has the following combination of favorable properties specific to PCCs. It is skewed along the first generator, which is the main common factor. The distribution is elliptical in the subspace of the higher generators, which are the orthogonal directions to the main factor resulting in diversification. The stochastic structure is based on two stochastic mixing variables. This differs from the commonly applied (skew) $t$ copula, which has a single stochastic mixing factor $W$. In that case, the stochastic variance factor $W$ grows for all directions simultaneously in extreme scenarios. The variation orthogonal to the parallel movement is not a priori expected to grow during a crash, because it enhances diversification. Consequently, we find in our empirical study of Section 4 that the Skew $t_{1}$-$t_{d-1}$ PCC specification leads to improvements in terms of information criteria and in capturing systemic risk compared to other copulas.